OpenStudy (anonymous):

The discriminant of a quadratic equation is the value b2 – 4ac. What is the value of the discriminant of 0 = –2x2 + 6x + 13?

OpenStudy (faiqraees):

general equation is \[\large\rm ax^2+bx+c=–2x^2 + 6x + 13 \] SO \[\large\rm b^2-4ac=6^2-4*(-2)*13 \]\[\large\rm discriminant=6^2-4*(-2)*13 \]